a vehicle leaves a point a with a speed of 5m.s and travels 20 m to reach 18m.s . Found the acceleration


1
Expert's answer
2021-09-23T08:30:48-0400
"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d}=\dfrac{(18\ \dfrac{m}{s})^2-(5\ \dfrac{m}{s})^2}{2\cdot20\ m}=7.47\ \dfrac{m}{s^2}."